### v1

@@ Abstract
Sequence models generate counterfactuals by modifying parts of a sequence based on a given condition, enabling reasoning about “what if” scenarios. While these models excel at conditional generation, they lack fine-grained control over when and where edits occur. Existing approaches either focus on univariate sequences or assume that interventions affect the entire sequence globally. However, many applications require precise, localized modifications, where interventions take effect only after a specified time and impact only a subset of co-occurring variables. We introduce Clef, a controllable sequence editing model for counterfactual reasoning about both immediate and delayed effects. Clef learns temporal concepts that encode how and when interventions should influence a sequence. With these concepts, Clef selectively edits relevant time steps while preserving unaffected portions of the sequence. We evaluate Clef on cellular and patient trajectory datasets, where gene regulation affects only certain genes at specific time steps, or medical interventions alter only a subset of lab measurements. Clef improves immediate sequence editing by up to $36.01\%$ in MAE compared to baselines. Unlike prior methods, Clef enables one-step generation of counterfactual sequences at any future time step, outperforming baselines by up to $65.71\%$ in MAE. A case study on patients with type 1 diabetes mellitus shows that Clef identifies clinical interventions that shift patient trajectories toward healthier outcomes.
@@ 1 Introduction
Figure 1: Illustrative comparison of (a) Clef’s controllable sequence editing and (b) existing sequence editing approaches. Unlike existing methods, controllable sequence editing generates counterfactual sequences (dotted lines) while preserving historical data to model the immediate or delayed effects of interventions.
Counterfactual thinking is a fundamental objective in biology and medicine Lee & Topol (2024). “What if” scenarios are critical for reasoning about the underlying mechanisms of a cell, patient, disease, and drug, and each decision can have tangible impact Bunne et al. (2024); Lee & Topol (2024): What if we treat the cells with the candidate drug every hour or every 24 hours? What if we perform the surgery on the patient today or next year? We should not only reason about the choice of the counterfactual condition (e.g., drug, surgery), but also its timing (e.g., when and how frequent). Thus, counterfactual generation requires precise and context-specific edits that adhere to temporal and structural constraints. For example, prescribing a medication to a patient should result in changes to the patient’s trajectory only after the intervention time (i.e., the medical history prior to intervention should be unaffected to preserve temporal causality) and on only the relevant variables that are specific to the context of the intervention (i.e., the measurements unaffected by the intervention should be preserved).
@@
Generative models in the language and vision domains enable precise editing guided by a description, such as textual prompts or condition tokens Zhang et al. (2023b); Gao et al. (2023); Ravi et al. (2024); Gong et al. (2024); Niu et al. (2024); Gu et al. (2024); Zhou et al. (2024). These models are designed to gain more comprehensive (i.e., global) and precise (i.e., local) control over the generation of text Chatzi et al. (2024); Niu et al. (2024); Gu et al. (2024); Zhou et al. (2024), images Zhang et al. (2023b); Gao et al. (2023); Ravi et al. (2024), and even molecular structures Gong et al. (2024); Dauparas et al. (2022); Zhang et al. (2024). Their outputs are expected to preserve the global integrity of the input while making precise local edits to satisfy the desired condition. Analogous to these models’ consideration of spatial context to edit images Zhang et al. (2023b); Gao et al. (2023) and protein pockets Dauparas et al. (2022); Zhang et al. (2024) via in-painting, our work leverages temporal context to perform precise editing on sequences.
@@
Controllable text generation (CTG) approaches, designed specifically to edit natural language sequences, have been extensively studied Zhang et al. (2023a). They excel in immediate sequence editing: predicting the next token or readout in the sequence under a counterfactual condition Niu et al. (2024); Gu et al. (2024); Zhou et al. (2024); Chatzi et al. (2024); Zhang et al. (2023a); Bhattacharjee et al. (2024). For example, if asked to predict the next word in the sentence “Once upon a time, there lived a boy” under the counterfactual condition that the genre is horror, a CTG model may respond with “alone” to convey vulnerability and loneliness. However, CTG models are unable to perform delayed sequence editing: predicting a counterfactual trajectory at a future time step while maintaining causal consistency. For example, if asked to predict what would happen to the boy if he took a trip to New York City as an adult, a CTG model would struggle due to the multitude of possible answers. In the existing paradigm of CTG models, they cannot effectively utilize the given context to skip ahead to the future; instead, CTG models would need to be run repeatedly to fill in the temporal gap without any guarantee of ever satisfying the desired condition. As a result, CTG models are insufficient for other types of sequences for which both immediate and delayed sequence editing are necessary, such as cellular reprogramming and patient immune dynamics.
@@
There exist two controllable time series generation approaches Jing et al. (2024); Bao et al. (2024), which utilize diffusion modeling to generate counterfactual time series. However, they are limited to univariate sequences and assume that the entire input sequence is affected Jing et al. (2024); Bao et al. (2024). These methods are thus insufficient in settings where edits are only allowed after time $t$ (i.e., cannot change historical data) and affect only certain sequences (i.e., preserve unaffected co-occurring sequences). In other words, they are unable to make precise local edits while preserving global causal consistency.
Present work. We tackle the gaps in controllable sequence editing to enable temporally localized modifications at any time step while ensuring the adherence to temporal causality and the consistency of intrinsic dependencies within and across sequences (Figure 1). Controllable sequence editing is a particularly complex task because it necessitates learning both the temporal dynamics in the sequences and the relationships between the desired condition and the sequences. The latter is most often unknown, hindering the model’s ability to determine (1) which sequences are affected by the condition, and (2) when and how the sequences are affected based on the historical data and the guiding condition. These challenges prevent sequence editing approaches from generating precise and context-specific edits while preserving temporal and structural constraints. Further, the resulting condition-guided counterfactual sequences may not resemble the distribution of observed sequences.
@@
We develop Clef (ControLlable sequence Editing for counterFactual generation), a controllable sequence editing approach for instance-wise counterfactual generation. Clef learns temporal concepts that represent the trajectories of the sequences to enable accurate counterfactual generation guided by a given condition. We show that the learned temporal concepts help preserve temporal and structural constraints in the generated outputs. By design, Clef is flexible with any type of sequential data encoder. We demonstrate through comprehensive experiments on four novel benchmark datasets in cellular reprogramming and patient immune dynamics that Clef outperforms state-of-the-art models by up to $36.01\%$ and $65.71\%$ (MAE) on immediate and delayed sequence editing, respectively. We also show that any pretrained sequence encoder can gain controllable sequence editing capabilities when finetuned with Clef. Moreover, Clef outperforms baselines in zero-shot counterfactual generation of cellular trajectories by up to $14.45\%$ and $63.19\%$ (MAE) on immediate and delayed sequence editing, respectively. Further, precise edits via user interaction can be performed directly on Clef’s learned concepts. We demonstrate through real-world case studies that Clef, given precise edits on specific temporal concepts, can generate realistic “healthy” counterfactual trajectories for patients originally with type 1 diabetes mellitus.
Our contributions are threefold. (1) We formalize controllable sequence editing for temporally localized modification of biomedical sequences at any time step given a desired condition while ensuring temporal causality. (2) Clef is a novel controllable sequence editing model for counterfactual generation. (3) We release four datasets on cellular reprogramming and patient immune dynamics, and demonstrate the efficacy of Clef on immediate and delayed sequence editing on cellular and patient trajectories.
@@ 2 Related work
Sequence editing. The sequence editing task has been defined in language and time series modeling via different terms, but share a core idea: Given a sequence and an alternative condition (e.g., sentiment, attribute), generate a counterfactual sequence with the desired properties. Counterfactual sequence generation is an autoregressive process in language Chatzi et al. (2024) but a diffusion process in time series Jing et al. (2024); Bao et al. (2024). Prompting is often used to guide the generation of a sequence, both textual and temporal, with a desired condition Zhang et al. (2023a); Bhattacharjee et al. (2024); Jing et al. (2024); Bao et al. (2024). However, existing approaches are unable to generate counterfactual multivariate sequences, preserve relevant historical data, and ensure time-sensitive interventions. Temporal sequence editing approaches assume that sequences are univariate and conditions affect the entire sequence Jing et al. (2024); Bao et al. (2024). While incorporating a structural causal model for token sampling can help preserve certain attributes during counterfactual text generation, the counterfactual statements may be inconsistent with real-world causal models Chatzi et al. (2024).
@@
Concept-based learning. Concepts can be thought of as abstract atomic ideas or concrete tokens of text or images The et al. (2024); Lai et al. (2024). Concept-based learning has been used to explain (e.g., predict the concepts observed in the sample) or transform black-box models into more explainable models (e.g., allow users to intervene on learned concepts) Koh et al. (2020); Shin et al. (2023); Ismail et al. (2024); Lai et al. (2024); Laguna et al. (2024); van Sprang et al. (2024). While concepts have been used in sequence generation, they have not yet been used for conditional generation. The adoption of concept-based learning for counterfactual prediction is limited to image classification, where concepts are intervened on during training to simultaneously learn the label and explanation Dominici et al. (2024). Further, there is a consistent and widely accepted trade-off between accuracy and interpretability in concept-based models.
Leveraging trajectories as inductive biases. Understanding sequential data as trajectories (e.g., increasing, decreasing, constant) is more natural for human interpretation than individual values Kacprzyk et al. (2024). Many modeling approaches on temporal data extract dynamic motifs as inductive biases to improve their interpretability Kacprzyk et al. (2024); Goswami et al. (2024); Cao et al. (2024). Such temporal patterns can be used for prompting large pretrained models to perform time series forecasting Cao et al. (2024), suggesting that trajectories can capture more universal and transferrable insights about the temporal dynamics in time series data. Trajectories have yet to be adopted for counterfactual sequence generation.
@@ 3 Clef
Clef manipulates sequences based on user-specified conditions and temporal coordinates to address “what if” questions. Given a sequence, a forecast time step, and a counterfactual condition, Clef modifies only the relevant portions of the sequence while preserving unaffected elements, ensuring causal consistency. For example, based on a patient’s historical lab test results, Clef generates a future lab test trajectory conditioned on a given treatment (Figure 1a). Architecturally, Clef has four key components: (i) a sequence encoder $F$ that extracts temporal features from historical sequence data, (ii) a condition adapter $H$ that maps counterfactual conditions to latent representations, (iii) a concept encoder $E$ that learns temporal concepts, representing trajectory patterns over time, and (iv) a concept decoder $G$ that applies these concepts to generate counterfactual sequences.
@@ 3.1 Problem definition
Figure 2: Overview of Clef’s architecture and capabilities. (a) Given an input sequence, forecast time, and condition embedding from a frozen pretrained (pt) embedding model, Clef generates a counterfactual sequence via immediate or delayed sequence editing. (b) Clef is composed of a sequence encoder, condition adapter, concept encoder, and concept decoder. Clef has two key capabilities: (c) forecasting counterfactual sequences at any time in the future and under any condition (e.g., medical codes), and (d) generating counterfactual sequences by intervening on Clef’s learned temporal concepts.
@@
Sequence editing is the local sample-level modification of sequence $\mathbf{x}$ to generate a counterfactual sequence $\hat{\mathbf{x}}_{:,t_{j}}$ under a given condition $s$ at a specific time $t_{j}$. There are two types of controllable sequence editing: immediate and delayed (Figure 2a).
  • Immediate sequence editing: Given sequence $\mathbf{x}_{:,t_{0}:t_{i}}$ and condition $s$ to occur at time $t_{i+1}$, forecast $\hat{\mathbf{x}}_{:,t_{i+1}}$.
  • Delayed sequence editing: Given sequence $\mathbf{x}_{:,t_{0}:t_{i}}$ and condition $s$ to occur at time $t_{j}\geq t_{i+1}$, forecast $\hat{\mathbf{x}}_{:,t_{j}}$.
Example scenarios for immediate sequence editing include: What if we perturb the cells now? and What if we perform surgery on the patient today? (Section 5.1). In contrast, delayed sequence editing applies to questions such as: What if we perturb the cells in ten days? and What if we perform surgery on the patient next year? (Section 5.2).
@@
A temporal concept $\mathbf{c}$ for sequence $\mathbf{x}_{:,t_{j}:t_{k}}$ is defined by $\mathbf{c}=\nicefrac{{\mathbf{x}_{:,t_{k}}}}{{\mathbf{x}_{:,t_{j}}}}$ for time steps $t_{j}$ and $t_{k}$ where $t_{k}>t_{j}$. It can be interpreted as the trajectory (or rate of change of each variable in the sequence) between any pair of time steps.
@@
Concept encoder $E$ and decoder $G$ are able to leverage temporal concepts $\mathbf{c}$ to perform controllable sequence editing on dataset $\mathcal{D}$ if the following are satisfied.
@@
  • Counterfactual condition $a\neq s$ on $\mathbf{x}_{:,t_{0}:t_{i}}$ at $t_{j}$ learns $\mathbf{c}^{\prime}\neq\mathbf{c}$ that forecasts $\hat{\mathbf{x}}^{a}_{:,t_{j}}$ such that $\hat{\mathbf{x}}^{a}_{:,t_{j}}\neq\hat{\mathbf{x}}^{s}_{:,t_{j}}$ and, if known, $\hat{\mathbf{x}}^{a}_{:,t_{j}}\simeq\mathbf{x}^{a}_{:,t_{j}}$.
@@
Given a sequence encoder $F$, condition adapter $H$, concept encoder $E$, and concept decoder $G$ trained on a longitudinal dataset $\mathcal{D}$, Clef learns temporal concept $\mathbf{c}=E(F(\mathbf{x}_{:,t_{0}:t_{i}},t_{j}),H(s))$ to forecast $\hat{\mathbf{x}}^{s}_{:,t_{j}}=G(\mathbf{x}_{:,t_{i}},\mathbf{c})$ for any sequence $\mathbf{x}_{:,t_{0}:t_{i}}\in\mathcal{D}$, future time step $t_{j}>t_{i}$, and condition $s$.
3.2 Clef model
The input to Clef are a multivariate sequence $\mathbf{x}_{:,t_{0}:t_{i}}$ with $V$ measured variables, and a condition $s$ and time $t_{j}>t_{i}$ for which to forecast $\hat{\mathbf{x}}^{s}_{:,t_{j}}$. Clef consists of four major components: a sequence encoder $F$, a condition adapter $H$, a concept encoder $E$, and a concept decoder $G$.
Sequence encoder $F$. The sequence encoder $F$ extracts features from $\mathbf{x}_{:,t_{0}:t_{i}}$ such that $\mathbf{h}_{\mathbf{x}}=F(\mathbf{x}_{:,t_{0}:t_{i}})$. Any encoder, including a pretrained multivariate foundation model, can be used. The time encoder in $F$ generates a time positional embedding $\mathbf{h}_{t}$ for any time $t$ via element-wise summation of the year (sinusoidal), month, date, and hour embeddings. It is additionally used to compute the time delta embedding $\Delta_{t_{i},t_{j}}=\mathbf{h}_{t_{j}}-\mathbf{h}_{t_{i}}$ for the concept encoder $E$.
Condition adapter $H$. The embedding $\mathbf{z}_{s}$ corresponding to the input condition $s$ is retrieved from a frozen pretrained embedding model (denoted as pt in Figure 2a). The condition adapter $H$ projects $\mathbf{z}_{s}$ into hidden representation $\mathbf{h}_{s}=H(\mathbf{z}_{s})$.

### v2

@@ Abstract
Conditional generation models for longitudinal sequences can generate new or modified trajectories given a conditioning input. While effective at generating entire sequences, these models typically lack control over the timing and scope of the edits. Most existing approaches either operate on univariate sequences or assume that the condition affects all variables and time steps. However, many scientific and clinical applications require more precise interventions, where a condition takes effect only after a specific time and influences only a subset of variables. We introduce Clef, a controllable sequence editing model for conditional generation of immediate and delayed effects in multivariate longitudinal sequences. Clef learns temporal concepts that encode how and when a condition alters future sequence evolution. These concepts allow Clef to apply targeted edits to the affected time steps and variables while preserving the rest of the sequence. We evaluate Clef on 7 datasets spanning cellular reprogramming and patient health trajectories, comparing against 9 state-of-the-art baselines. Clef improves immediate sequence editing accuracy by up to $36.01\%$ (MAE). Unlike prior models, Clef enables one-step conditional generation at arbitrary future times, outperforming them in delayed sequence editing by up to $65.71\%$ (MAE). We test Clef under counterfactual inference assumptions and show up to $63.19\%$ (MAE) improvement on zero-shot conditional generation of counterfactual trajectories. In a case study of patients with type 1 diabetes mellitus, Clef identifies clinical interventions that generate realistic counterfactual trajectories shifted toward healthier outcomes.
@@ 1 Introduction
Figure 1: Illustrative comparison of (a) controllable sequence editing and (b) existing sequence editing. Unlike existing methods, controllable sequence editing generates sequences (dotted lines) guided by a condition while preserving historical data to model the effects of immediate (e.g., in 2 hours) or delayed (e.g., in 1 week) edits.
As the US Food and Drug Administration (FDA) phases out animal testing requirements for therapeutic discovery, AI tools will become widely adopted to simulate the effects of candidate drugs [fda]. In particular, building virtual cells and patients to model the behavior of molecules, cells, and tissues could facilitate large-scale in silico experimentation, such as approximating the efficacy of a drug [bunne2024build, li2025digital]. In this work, we focus on the conditional generation of longitudinal sequences. We posit that conditional sequence generation models can be used to infer potential outcomes under specific conditions: Generate a cell’s state after treating it with a candidate drug every hour or every 24 hours. Generate a patient’s state after performing surgery today or next year. To successfully build virtual cells and patients, we must be able to reason about both the choice of the intervention (e.g., drug, surgery) and its timing (e.g., when and how frequent). Thus, conditional generation on longitudinal sequences requires precise and context-specific edits. For example, prescribing a medication to a patient should result in changes to the patient’s trajectory only after the intervention time (i.e., the medical history prior to intervention should be unaffected) and on only the relevant variables that are specific to the context of the intervention (i.e., the measurements unaffected by the intervention should be preserved).
@@
Generative language and vision models enable precise editing guided by a description, such as textual prompts or condition tokens [zhang2023adding, gao2023editanything, ravi2024sam2, gong2024text, niu2024mitigating, gu2024token, zhou2024virtual]. These models are designed to gain more global context-preserving and local precise control over the generation of text [chatzi2024counterfactual, niu2024mitigating, gu2024token, zhou2024virtual], images [zhang2023adding, gao2023editanything, ravi2024sam2], and even molecular structures [gong2024text, dauparas2022robust, zhang2024efficient]. Their outputs preserve the input’s global integrity yet contain precise local edits to satisfy the desired condition. Analogous to these models’ consideration of spatial context to edit images [zhang2023adding, gao2023editanything] and protein pockets [dauparas2022robust, zhang2024efficient] via in-painting, our work leverages temporal context to edit sequences based on a given condition.
@@
Controllable text generation (CTG) approaches, designed specifically to edit natural language sequences, have been extensively studied [zhang2023survey]. They excel in immediate sequence editing: predicting the next token or readout in the sequence under a given condition [niu2024mitigating, gu2024token, zhou2024virtual, chatzi2024counterfactual, zhang2023survey, bhattacharjee2024zero]. For example, if asked to predict the next word in the sentence “Once upon a time, there lived a boy" under the condition that the genre is horror, a CTG model may respond with “alone" to convey vulnerability and loneliness. However, CTG models are unable to perform delayed sequence editing: modifying a trajectory in the far-future. The distinction is important: the focus is on when the edit occurs, not necessarily when its effects manifest. Whereas immediate sequence editing applies a condition now (e.g., administering insulin today), delayed sequence editing schedules a condition for the future (e.g., starting a chemotherapy regimen in six weeks). Existing CTG models cannot effectively utilize the given context to skip ahead to the future; instead, they would need to be run repeatedly to fill in the temporal gap without any guarantee of satisfying the desired condition. As a result, CTG models are insufficient for other sequence types (i.e., not natural language) for which both immediate and delayed sequence editing are necessary, such as cell development and patient health trajectories.
@@
There exist two controllable time series generation (CTsG) approaches [jing2024towards, bao2024towards], which utilize diffusion modeling to generate time series under a given condition. However, they are limited to univariate sequences and assume that the entire input sequence is affected [jing2024towards, bao2024towards]. These methods are thus insufficient in settings where edits are only allowed after time $t$ (i.e., cannot change historical data) and affect only certain sequences (i.e., preserve unaffected co-occurring sequences). In other words, CTsG methods are unable to make precise local edits while preserving global integrity. Orthogonal to CTsG is the estimation of counterfactual outcomes over time (ECT) [melnychuk2022causal, bica2020estimating, huang2024empirical, wang2025effective]. Although not generative, ECT autoregressively predicts the potential outcomes (i.e., next readout in the sequence) as a result of different future treatments (i.e., fixed set of conditions) under counterfactual inference assumptions. While ECT preserves historical and unaffected co-occurring sequences, counterfactual inference assumptions may not always hold in real-world applications.
@@
Present work. We develop Clef  [^1] GitHub repository: https://github.com/mims-harvard/CLEF, a novel ControLlable sequence Editing Framework for instance-wise conditional generation. Clef learns temporal concepts that represent the trajectories of the sequences to enable accurate generation guided by a given condition (Definition 3.2). We show that the learned temporal concepts help preserve temporal constraints in the generated outputs. By design, Clef is flexible with any type of sequential data encoder. We demonstrate through experiments on 4 new benchmark datasets in cellular reprogramming and patient immune dynamics that Clef outperforms state-of-the-art models by up to $36.01\%$ and $65.71\%$ (MAE) on immediate and delayed sequence editing (Definition 3.1). We also show that any pretrained sequence encoder can gain controllable sequence editing capabilities when finetuned with Clef. Additionally, we extend Clef for multi-step ahead counterfactual prediction under counterfactual inference assumptions (Assumption 3.4, Equation 5), and demonstrate (on 3 benchmark datasets) performance gains against 5 state-of-the-art baselines in settings with high time-varying confounding. Moreover, Clef enables conditional generation models to outperform baselines in zero-shot generation of counterfactual cellular trajectories by up to $14.45\%$ and $63.19\%$ (MAE) on immediate and delayed sequence editing. Further, precise edits via user interaction can be performed directly on Clef’s learned concepts. We show through real-world case studies that Clef, given precise edits on specific temporal concepts, can generate realistic “healthy" trajectories for patients originally with type 1 diabetes mellitus.
Our contributions are fourfold. (1) We develop Clef: a flexible controllable sequence editing model for conditional generation of longitudinal sequences. (2) Clef can be integrated into the (balanced) representation learning architectures of counterfactual prediction models to estimate counterfactual outcomes over time. (3) Beyond achieving state-of-the-art performance in conditional sequence generation and counterfactual outcomes prediction, Clef excels in zero-shot conditional generation of counterfactual sequences. (4) We release four new benchmark datasets on cell reprogramming and patient immune dynamics for immediate and delayed sequence editing, and evaluate on three established benchmark datasets regarding patient tumor growth for counterfactual prediction.
@@ 2 Related work
Sequence editing. The sequence editing task has been defined in language and time series modeling via different terms, but share a core idea: Given a sequence and a condition (e.g., sentiment, attribute), generate a sequence with the desired properties. Conditional sequence generation is an autoregressive process in language [chatzi2024counterfactual] but a diffusion process in time series [jing2024towards, bao2024towards]. Prompting is often used to guide the generation of a sequence, both textual and temporal, with a desired condition [zhang2023survey, bhattacharjee2024zero, jing2024towards, bao2024towards]. However, existing approaches are unable to generate multivariate sequences, preserve relevant historical data, and ensure time-sensitive edits. They assume that sequences are univariate and conditions affect the entire sequence [jing2024towards, bao2024towards]. Structural causal models can be incorporated to enable counterfactual text generation while preserving certain attributes [chatzi2024counterfactual, ravfogel2025gumbel]. Estimating counterfactual outcomes over time is often formulated under the potential outcomes framework [neyman1923application, rubin1978bayesian].
Estimating counterfactual outcomes over time. Predicting time-varying counterfactual outcomes in time series entails estimating counterfactual outcomes over possible sequences of interventions, including the timing and ordering of sequential treatments [melnychuk2022causal, bica2020estimating, huang2024empirical, wang2025effective]. There are decades of research on temporal counterfactual outcomes estimation [robins2000marginal, lim2018forecasting, bica2020estimating, li2021gnet, melnychuk2022causal]. Recently, machine learning approaches for predicting time-varying counterfactual outcomes learn representations that are predictive of outcomes while mitigating treatment bias via balancing techniques [melnychuk2022causal, bica2020estimating, huang2024empirical, wang2025effective]. On image data, conditional generation models (i.e, guided diffusion, conditional variational autoencoder) have been shown to predict counterfactual outcomes without an explicit density estimation [wu2024counterfactual]. However, there may be a trade-off between prediction accuracy and balanced representations [huang2024empirical].
@@
Concept-based learning. Concepts can be thought of as abstract atomic ideas or concrete tokens of text or images [the2024large, lai2023faithful]. Concept-based learning has been used to explain (e.g., predict the concepts observed in the sample) or transform black-box models into more explainable models (e.g., allow users to intervene on learned concepts) [koh2020concept, shin2023closer, ismail2023concept, lai2023faithful, laguna2024beyond, van2024enforcing]. While concepts have been used in sequence generation [the2024large], they have not yet been used for conditional generation of longitudinal sequences. The adoption of concept-based learning for counterfactual prediction is limited to improving the interpretability of image classification [dominici2024climbing, de2025causally, dominici2024causal].
Refer to Appendix A for further discussion of related work.
@@ 3 Clef
Clef manipulates sequences based on user-specified conditions and temporal coordinates. Given a longitudinal sequence, forecast time step, and condition token, Clef modifies only the relevant portions of the sequence while preserving unaffected elements, ensuring global integrity. Architecturally, Clef has four key components: (i) sequence encoder $F$ that extracts temporal features from historical sequence data, (ii) condition adapter $H$ that maps condition tokens to latent representations, (iii) concept encoder $E$ that learns temporal concepts, representing trajectory patterns over time, and (iv) concept decoder $G$ that applies these concepts to generate sequences.
Figure 2: Overview of Clef’s architecture and capabilities. (a) Given a sequence, forecast time, and condition embedding from a frozen pretrained (pt) embedding model, Clef generates a sequence via immediate or delayed sequence editing. (b) Clef is composed of a sequence encoder, condition adapter, concept encoder, and concept decoder. Clef has two key capabilities: (c) forecast sequences at any future time and under any condition (e.g., medical codes), and (d) generate sequences by intervening on Clef’s learned temporal concepts.
@@ 3.1 Problem definition
Consider an observational dataset $\mathcal{D}=\{\mathbf{x}^{(i)}_{t},\mathbf{s}^{(i)}_{t}\}^{N}_{i=1}$ for $N$ independent entities (e.g., cells, patients) at time step $t$. For each entity $i$ at time $t$, we observe continuous time-varying covariates $\mathbf{x}^{(i)}_{t}\in\mathbb{R}^{d_{x}}$ (e.g., gene expression, laboratory test measurements) and categorical conditions $\mathbf{s}^{(i)}_{t}$ (e.g., transcription factor activation, clinical intervention). The outcome of the condition is measured by the covariates (e.g., a transcription factor activation affects a cell’s gene expression, a medication affects a patient’s laboratory test profile). For notation, we omit entity index $(i)$ unless needed.
@@
Sequence editing is the local sample-level modification of sequence $\mathbf{x}$ to autoregressively generate $\hat{\mathbf{x}}_{:,t_{j}}$ under condition $s$ given at time $t_{j}-\epsilon$. Time gap $\epsilon$ indicates that $\hat{\mathbf{x}}_{:,t_{j}}$ is measured a negligible amount of time after $s$ is applied [^2] We assume that the condition $s$ always occurs shortly before the measured covariates at time $t_{j}$. When $s$ is an intervention and our problem becomes counterfactual prediction (refer to Section 3.3 for a more rigorous discussion about the formulation of sequence editing in this context), our assumption is consistent with existing practice in the literature of counterfactual prediction.; for notation, we omit $\epsilon$ unless needed. There are two types of controllable sequence editing (Figure 2a):
  • Immediate sequence editing: Given $\mathbf{x}_{:,t_{0}:t_{i}}$ and $s$ to occur at $t_{i+1}$, forecast $\hat{\mathbf{x}}_{:,t_{i+1}}$
  • Delayed sequence editing: Given $\mathbf{x}_{:,t_{0}:t_{i}}$ and $s$ to occur at $t_{j}\geq t_{i+1}$, forecast $\hat{\mathbf{x}}_{:,t_{j}}$
Examples of immediate sequence editing include generating trajectories after perturbing cells now or performing surgery on patients today (Section 5.1). In contrast, delayed sequence editing generates trajectories after perturbing cells in ten days or performing surgery on patients next year (Section 5.2).
@@
Temporal concept $\mathbf{c}$ is defined by $\mathbf{c}=\nicefrac{{\mathbf{x}_{:,t_{k}}}}{{\mathbf{x}_{:,t_{j}}}}$ for sequence $\mathbf{x}$ where time steps $t_{k}>t_{j}$. It can be interpreted as the trajectory (or rate of change of each variable in the sequence) between any pair of time steps.
@@
Concept encoder $E$ and decoder $G$ can leverage temporal concepts $\mathbf{c}$ to perform controllable sequence editing if the following are satisfied.
@@
  • For an alternative condition $a\neq s$ on $\mathbf{x}_{:,t_{0}:t_{i}}$ at $t_{j}$, the method learns a distinct $\mathbf{c}^{\prime}\neq\mathbf{c}$ that forecasts $\hat{\mathbf{x}}^{a}_{:,t_{j}}$ such that $\hat{\mathbf{x}}^{a}_{:,t_{j}}\neq\hat{\mathbf{x}}^{s}_{:,t_{j}}$ and, if known, $\hat{\mathbf{x}}^{a}_{:,t_{j}}\simeq\mathbf{x}^{a}_{:,t_{j}}$.
@@
Given a sequence encoder $F$, condition adapter $H$, concept encoder $E$, and concept decoder $G$ trained on a longitudinal dataset $\mathcal{D}$, Clef learns temporal concept $\mathbf{c}=E(F(\mathbf{x}_{:,t_{0}:t_{i}},t_{j}),H(s))$ to forecast $\hat{\mathbf{x}}^{s}_{:,t_{j}}=G(\mathbf{x}_{:,t_{i}},\mathbf{c})$ for any $\mathbf{x}_{:,t_{0}:t_{i}}\in\mathcal{D}$, $t_{j}>t_{i}$, and $s$.
$\hat{\mathbf{x}}^{s}_{:,t_{j}}=G(\mathbf{x}_{:,t_{i}},E(F(\mathbf{x}_{:,t_{0}:t_{i}},t_{j}),H(s)))$ (1)
3.2 Clef architecture
The input to Clef is a continuous multivariate sequence $\mathbf{x}_{:,t_{0}:t_{i}}\in\mathbb{R}^{V}$ with $V$ measured variables, a condition $s$, and time $t_{j}>t_{i}$ for which to forecast $\hat{\mathbf{x}}^{s}_{:,t_{j}}$. Clef consists of four major components: a sequence encoder $F$, a condition adapter $H$, a concept encoder $E$, and a concept decoder $G$.
Sequence encoder $F$. The sequence encoder $F$ extracts features $\mathbf{x}_{:,t_{0}:t_{i}}$ such that $\mathbf{h}_{\mathbf{x}}=F(\mathbf{x}_{:,t_{0}:t_{i}})$. Any encoder, including a pretrained multivariate foundation model, can be used. The time encoder in $F$ generates a time positional embedding $\mathbf{h}_{t}$ for any time $t$ via element-wise summation of the year (sinusoidal), month, date, and hour embeddings. It is also used to compute the time delta embedding $\Delta_{t_{i},t_{j}}=\mathbf{h}_{t_{j}}-\mathbf{h}_{t_{i}}$ for the concept encoder $E$.
Condition adapter $H$. The condition token, or embedding $\mathbf{z}_{s}$ corresponding to the input condition $s$, is retrieved from a frozen pretrained embedding model (denoted as pt in Figure 2a). The condition adapter $H$ projects $\mathbf{z}_{s}$ into a hidden representation $\mathbf{h}_{s}=H(s)$.